Tohban Report 2012-12-26
From RHESSI Wiki
|Start Date:||19 Dec 2012|
|End Date:||26 Dec 2012|
|Next Tohban:||Pascal Saint Hilaire (firstname.lastname@example.org)|
|List all reports|
The Sun remained quiet this week, but solar activity is now picking back up as one of 4 currently active regions on the disk (AR 11635) has produced several C flares during the last two days. STEREO B shows one further active region which will rotate onto the disk in the next few days.
How many GOES flares occurred?
Flares above B, C, M, X class were 32 16 0 0
And how many of these are listed in the RHESSI flare list?
Flares above B, C, M, X class were 24 11 0 0
And how many had EXCELLENT coverage?
Flares above B, C, M, X class were 13 9 0 0
There were RHESSI flares/GOES flares 200 / 48 over the time range 19-Dec-12 26-Dec-12
The SSR fill level was relatively high earlier this week because of replaying data in response to the ground station problems from the previous week.
The CPC temperature was high but trending normally with day length.
There was a large data gap due to the Wallop station problem.
GAPS IN APP_ID = 102 (VC3-MONITOR RATES) WITH PACKET RATE LT 30 GAP START TIME GAP END TIME GAP (SEC) 2012-12-20T01:10:00.000 -- 2012-12-20T04:55:00.000 13500.000
Here is the high voltage changes made this week.
2012-356-20:07:08 /ihvdac detector=8, voltage=127 (2475.5 V) 2012-356-20:08:51 /ihvdac detector=4, voltage=221 (4313.8 V)
Detector 8 seems to have been quiet since this change. It looks like it actually quieted before the HV adjustment, although some of the data from that interval still shows noise. Detector 8 should be turned back on next week by the Ops team at their earliest convenience.
As for Detector 4, we don't see signs of desegmenting in the most recent plots, other than very briefly during the SAA.
|HLAT Decimation||Rear decimation weight 6, no front decimation|
|Night time data (fronts)||+/- 4 min|
|Night time data (rears)||+/- 4 min|
|Require extra passes?||Yes|
|Requirement for moving pointer?||Some data replayed|
|Detector problems?||D4 and D8|